Four hundred pieces of small ice (m = 3 g) are poured into a Thermos of water in preparation for a soccer practice. There are 3 gallons of water. The water is at 70 ˚F to start. The system has enough time to reach thermal equilibrium before being brought out to practice. Will the water taste refreshing to the players? The water temperature must be 60 ˚F or below to be perceived as "refreshing."
Answer #1

Water will taste refreshing to the players.

Heat capacity of water= 42*10 35/K9C Latent heat of fusion of icl = 3.35 x 155/kg Water Mw = 3*3-785 = 1.355 kg Ice Me = 400*